OpenStudy (anonymous):

solve for x: 5x+2=27 solve for x: 5x+2=27 @Mathematics

OpenStudy (anonymous):

5x=25 divide by 5 x=5 check by plugging back in

OpenStudy (anonymous):

x = 5

OpenStudy (anonymous):

try 5

OpenStudy (anonymous):

5x+2=27 -2 -2 5x=25 5/5 = 25/5 x=5

OpenStudy (anonymous):

First step: get ALL xs on one side, get everything on the other. then try and get x by itself. That's the logic behind it.